Sol Summary Report – February 9th

Crew 188 Sol 12 Summary Report 9FEB2018

Sol 12

Summary Title: One more round then it’s [oxygen] bottles to the ground

Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander

Mission Status: EVAs are complete, it’s the final crunch.

Sol Activity Summary:

As the Olympic flame is lit back on Earth, our physical challenges in the field come to a close. With two final EVAs, the crew is dusting off the suits and preparing them for the next crew who will start a week later. The morning EVA hiked up to the Hab Ridge for some final 360 filming, chlorophyll sampling and scenic views. The afternoon EVA had a few samples along the way to Toothy Ridge to the North. Final working photos were snapped with a few last planned for tomorrow during our campus cleanup.

As with any online event on Earth, our Mars-2-Mars event will have a slight change of plans with the session being recorded on WebEx and shared immediately afterwards by the Austrian Space Forum (Österreichisches Weltraum Forum, OeWF). Q&A can be submitted to reddit and the crews will address them during the hangout or afterwards as an ongoing discussion.

10Feb2018, 17:00 CET = 09:00 MDRS time (Sol 13) = 20:00 Oman time [16:00 GMT]

reddit Q&A: https://www.reddit.com/r/Mars/comments/7s23yz/mdrs_crew_188_ama

Spacesuit Up!

Ryan L. Kobrick, Ph.D.

MDRS Crew 188 Commander

Look Ahead Plan:

A hangout with #SimulateMars in Oman, the AMADEE-18 mission at Kepler Station. Then we will end the MDRS Crew 188 simulation mid-day as we continue to clean all areas. With our continuous Hab-chores cleaning schedule we should be on track for an early Sunday morning departure.

Anomalies in work:

None.

Weather:

Calm winds and hazy skies with warm dropping all day.

Crew Physical Status:

There is a mixed of anxiety to get work completed, excitement to return home to loved ones, and sadness to be separating form our Martian family.

EVA:

15 EVAs were completed during our mission.

Reports to file:

1. Ops Report

2. Sol Summary

3. Journalist Report

4. Green Hab Report

5. EVA #14 and #15 report

6. Daily Photos

7. Health & Safety Officer Report

8. Mission Summary

Support Requested (same as previous three reports):

– Heater for science dome requested. A purchase order request was made.

– FYI for mission support to be aware of for systems to track and safe operations.

– There are potentially two packages at Hollow Mountain that we would love to have picked up.

– We will pick them up post-sim

– Vegetables were requested due to the insufficient supply for six crewmembers. Non-vegetarians have cut their personal rations to help stretch the supplies for the two vegetarians. We have two celiac members who cannot eat the starches provided. These diets were all communicated before our arrival. In order to feed crewmembers over the next 4 days, 2-3 cans of dried vegetables were requested (or fresh) with the missing items from the original food supply. We were informed today that we will not be getting any form of resupply. Susan Holden Martin suggested we reiterate this request to mission support since it remains open.